Use the Laplace transform to solve the given initial-value problem.


y′ + 2y = sin 4t, y(0) = 1


1
Expert's answer
2022-05-19T14:52:40-0400

the Laplace transform for the first derivative:


"L\{y'\}=sY(s)-y(0)"

then for IVP:


"L\\{y'\\}+2 L\\{y\\}=L\\{sin\u20614t\\}""sY(s)-y(0)+2Y(s)=\frac 4 {s^2+16}""Y(s)(s+2)=\frac 4 {s^2+16}+1""Y(s)=\frac 4 {(s^2+16)(s+2)}+\frac 1 {s+2}""=\frac {s^2+20} {(s^2+16)(s+2)}=\frac {As+B} {s^2+16}+\frac C {s+2}""s^2+20=As^2+Bs+2As+2B+Cs^2+16C""s^2+20=(A+C)s^2+(2A+B)s+2B+16C""\begin{cases} A+C=1 \\ 2A+B=0 \\ 2B+16C=20 \end{cases} \rightarrow A=-0.2, B=0.4, C=1.2""Y(s)=\frac {-0.2s+0.4} {s^2+16}+\frac {1.2} {s+2}""y(t)=L^{-1} \{Y(s)\}=L^{-1}\bigg\{\frac{-0.2s+0.4}{s^2+16}\bigg\}+L^{-1}\bigg\{\frac{1.2}{s+2}\bigg\}""=-0.2L^{-1}\bigg\{\frac{s}{s^2+16}\bigg\}+\frac{0.4}{4}L^{-1}\bigg\{\frac{4}{s^2+16}\bigg\}+1.2L^{-1}\bigg\{\frac{1}{s+2}\bigg\}""=-0.2 cos(\u20614t)+0.1 sin\u2061(4t)+1.2 e^{-2t}"

Answer: "y(t)=-0.2 cos(\u20614t)+0.1 sin\u2061(4t)+1.2 e^{-2t}"